OpenStudy (anonymous):

The conservation club washed 24 cars and earned $144 on Saturday. At this rate, how much money did they make on Sunday washing 30 cars? A. $198 B. $180 C. $174 D. $150

OpenStudy (anonymous):

I will medal!

iYuko (iyuko):

144 / 24 = 6. 6 x 30 = 180. Therefore washing 30 cars will be $180

iYuko (iyuko):

So B

OpenStudy (texaschic101):

144/24 = x/30 --- ($144 to 24 cars = $x to 30 cars) this is a proportion, so we must now cross multiply (24)(x) = (144)(30) 24x = 4320 x = 4320/24 x = 180 so they earned $180 washing 30 cars

OpenStudy (anonymous):

Thank you Tex

iYuko (iyuko):

I said the same thing he said...Why he get the medal?